OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -2 and 16, respectively

OpenStudy (raffle_snaffle):

Can you please show me the basic geometric function?

OpenStudy (anonymous):

Im honestly so lost with this one

OpenStudy (anonymous):

see the fith term is given and the second term is given then we can find the common ratio..

OpenStudy (anonymous):

to know a geometric sequence we need to know the first term and the common ratio so if a be the first term and r be the common ratio then the nth term of the geometric sequence is given by a* r^(n-1)

OpenStudy (anonymous):

thus second term is given by ar= -2 and the fifth term is given by ar^4 = 16

OpenStudy (anonymous):

so using the above eq. we need to know a and r here first..

OpenStudy (anonymous):

thus ar^4 /ar =16/-2 or r^3= -8 =(-2)^3 hence r=-2 now ar =-2 so a= 1

OpenStudy (anonymous):

do i start putting it into the fromula

OpenStudy (anonymous):

yup as u wish.. thus nth term is given by = 1 *(-2)^(n-1)...

OpenStudy (anonymous):

Thank you so much!